### v1

@@ Abstract
Clinicians, as well as dermatology and pathology trainees, often struggle to access high-quality, open-access dermatopathology image datasets for diagnostic or educational purposes 1, 2, 3, 4, 5, 6. Physical or digital slides used for teaching are typically proprietary to a specific institution and may include patient identifiers, restricting their broader use7, 8, 9. Some sites, such as VisualDx, offer a limited collection of dermatopathology images but require a user subscription10. Others, like Pathologyoutline.com, provide only a small number of images illustrating classic disease features11. While PathPresenter offers high-resolution digital collections of scanned slides that can be uploaded by any user, not all cases on the platform are peer-reviewed, which may affect reliability12.
In this study, we developed a pipeline to accelerate the automated construction of a dermatopathology image dataset using the PubMed Central Open Access Subset (PMC-OA) as our source13. PMC-OA is a free, peer-reviewed, open-access platform that hosts an extensive range of biomedical information14. Specifically, we curated and cataloged images from PMC-OA that correspond to common benign and malignant cutaneous neoplasms, aiming to create a comprehensive dermatopathology dataset for educational purposes, cross-referencing, and machine learning applications. Our approach leveraged a deep learning (DL)-enabled algorithm to facilitate the identification and organization of relevant images. Additionally, we evaluated and compared the precision and recall metrics of traditional keyword-based retrieval against those of our DL-based retrieval algorithms.
@@ Methods
Using the PubMed API (Entrez Programming Utilities15), we then searched for PMC-OA articles containing relevant keywords in their titles or abstracts. For instance, the query term for identifying articles related to Angiokeratoma was “Angiokeratoma [Title/Abstract].” The E-utilities tool retrieved the PubMed Central ID (PMCID) for each relevant article, which serves as the unique reference number within the PMC-OA repository.
Next, we used these PMCIDs to obtain the full-text articles in XML format, which adheres to the NISO standard DTD for article publishing16. We then parsed the XML files to extract figures and their corresponding legends. When figures were presented as compounds containing multiple panels, we separated them into individual panels using an existing deep-learning model developed by Yao et al17.
@@
First, we trained a deep learning-based image modality classifier using the ImageCLEF 2016 dataset (Table 1)18. Our model architecture was based on DenseNet-121, pre-trained on ImageNet19. We replaced the original classification layer with a fully connected layer followed by a sigmoid activation function, enabling the model to output the probability of an input image being a pathology image. All images were resized to 224 $\times$ 224 $\times$ 3 to standardize input dimensions.
@@
The DermpathNet dataset is distributed as a compressed archive containing images (images.zip), along with three structured metadata files 20. All images are consistently named according to the filename field recorded in the metadata. The file metadata_figures.csv links each image to its source article and disease annotation, containing information such as the PubMed Central identifier (pmcid), the associated disease label, the filename, and the original figure caption. The file metadata_ontology.csv provides standardized disease information using controlled vocabularies, including disease, category, type, UMLS, preferred Name, MESH, and synonyms. The file metadata_articles.csv documents the article-level attributes from which the images were derived, including the pmcid and pmid, article title, publishing journal, and license information. Together, these resources allow users to explore DermpathNet at the level of individual images, disease categories, or source articles, while also ensuring transparent documentation of licensing for all included content.

### v2

@@ Abstract
Clinicians, as well as dermatology and pathology trainees, often struggle to access high-quality, open-access dermatopathology image datasets for diagnostic or educational purposes 15, 18, 8, 12, 6, 17. Physical or digital slides used for teaching are typically proprietary to a specific institution and may include patient identifiers, restricting their broader use2, 7, 9. Some sites, such as VisualDx, offer a limited collection of dermatopathology images but require a user subscription16. Others, like Pathologyoutline.com, provide only a small number of images illustrating classic disease features11. While PathPresenter offers high-resolution digital collections of scanned slides that can be uploaded by any user, not all cases on the platform are peer-reviewed, which may affect reliability14.
In this study, we developed a pipeline to accelerate the automated construction of a dermatopathology image dataset using the PubMed Central Open Access Subset (PMC-OA) as our source3. PMC-OA is a free, peer-reviewed, open-access platform that hosts an extensive range of biomedical information10. Specifically, we curated and cataloged images from PMC-OA that correspond to common benign and malignant cutaneous neoplasms, aiming to create a comprehensive dermatopathology dataset for educational purposes, cross-referencing, and machine learning applications. Our approach leveraged a deep learning (DL)-enabled algorithm to facilitate the identification and organization of relevant images. Additionally, we evaluated and compared the precision and recall metrics of traditional keyword-based retrieval against those of our DL-based retrieval algorithms.
@@ Methods
Using the PubMed API (Entrez Programming Utilities13), we then searched for PMC-OA articles containing relevant keywords in their titles or abstracts. For instance, the query term for identifying articles related to Angiokeratoma was “Angiokeratoma [Title/Abstract].” The E-utilities tool retrieved the PubMed Central ID (PMCID) for each relevant article, which serves as the unique reference number within the PMC-OA repository.
Next, we used these PMCIDs to obtain the full-text articles in XML format, which adheres to the NISO standard DTD for article publishing1. We then parsed the XML files to extract figures and their corresponding legends. When figures were presented as compounds containing multiple panels, we separated them into individual panels using an existing deep-learning model developed by Yao et al20.
@@
First, we trained a deep learning-based image modality classifier using the ImageCLEF 2016 dataset (Table 1)4. Our model architecture was based on DenseNet-121, pre-trained on ImageNet5. We replaced the original classification layer with a fully connected layer followed by a sigmoid activation function, enabling the model to output the probability of an input image being a pathology image. All images were resized to 224 $\times$ 224 $\times$ 3 to standardize input dimensions.
@@
The DermpathNet dataset is distributed as a compressed archive containing images (images.zip), along with three structured metadata files 19. All images are consistently named according to the filename field recorded in the metadata. The file metadata_figures.csv links each image to its source article and disease annotation, containing information such as the PubMed Central identifier (pmcid), the associated disease label, the filename, and the original figure caption. The file metadata_ontology.csv provides standardized disease information using controlled vocabularies, including disease, category, type, UMLS, preferred Name, MESH, and synonyms. The file metadata_articles.csv documents the article-level attributes from which the images were derived, including the pmcid and pmid, article title, publishing journal, and license information. Together, these resources allow users to explore DermpathNet at the level of individual images, disease categories, or source articles, while also ensuring transparent documentation of licensing for all included content.